# These authors contributed equally to this work

### v1

@@ Abstract
Analog bolometers based on temperature-dependent phase-transition materials such as vanadium oxide (VOx) and barium titanate (BTO) represent the state of the art in uncooled infrared detectors. Recently, the first room-temperature spintronic Poisson bolometer based on magnetic tunnel junctions (MTJs) was proposed and demonstrated as a promising infrared detector. Unlike conventional bolometers, the spintronic Poisson bolometer operates in a probabilistic regime dominated by Poissonian noise, where the response is governed by resistance fluctuations arising from thermally activated magnetization transitions. Spontaneous transitions between two metastable magnetic states occur even in the absence of incident light, and the transition probability increases under illumination. In this work, we experimentally study the statistical properties of the optical response of the spintronic Poisson bolometer under illumination. We demonstrate that transitions in spintronic Poisson bolometers, both in the absence and presence of light, exhibit Poissonian behavior, with transition rates and interarrival times modulated by incident radiation. Under illumination, we observe a 153% increase in the count rate accompanied by a 70% reduction in interarrival time. These results establish spintronic Poisson bolometers as a promising platform for probabilistic, high-speed, and high-sensitivity infrared detection at room temperature.
@@ I Introduction
Bolometers are thermal detectors in which incident light causes a temperature-induced resistance change that is read out electrically YADAV2022113611. Infrared bolometers are essential for emerging applications in thermal imaging, autonomous navigation, astronomy, and time-of-flight sensing Bao2023; Lahiri2012; Sivaprakasam2020; Kaushal2017; anderson2023advancements; rodriguez2023remote; PhysRevX.14.041005. State-of-the-art uncooled bolometers are analog detectors based on temperature-dependent phase transition of materials such as vanadium oxide (VOx) and barium titanate (BTO) yeh2020performance; scott2022sensing. Other uncooled bolometers are based on semiconductor and metallic materials with temperature-dependent resistance, including amorphous silicon and platinumdoi:10.1021/acsphotonics.9b01198; doi:10.1021/acs.nanolett.1c02972; doi:10.1021/acs.nanolett.3c03076. In VOx bolometers, major noise sources are Johnson noise, thermal noise, generation–recombination and $1/f$ noise philipose2023analyzing; 10.1063/5.0049633; du2013graphene. Similarly, the dominant noise in barium titanate (BTO) bolometers and graphene bolometers is Johnson noise Hanel:61; du2013graphene; PhysRevB.96.165431. All of the uncooled bolometers aforementioned are analog detectors that produce a continuous voltage or current proportional to the incoming signal, and are dominated by Gaussian noise, as shown in Fig. 1(a).
@@
Recently, the first uncooled spintronic Poisson bolometer, called the spintronic ultrafast nanoscale bolometer (SUN bolometer) leif_nanolett was proposed. Unlike analog bolometers, the spintronic Poisson bolometers convert incoming signals into discrete counting events and are dominated by Poissonian noise, as shown in Fig. 1(b). In spintronic Poisson bolometers, magnetic tunnel junctions (MTJs) are engineered with an energy barrier of the order of thermal infrared energy at room temperature. It exploits switching between two discrete magnetization states by intrinsic thermal fluctuations and nanoscale heating induced by incident radiation. These digital probabilistic transitions in the absence of light are referred to as dark counts, and when under illumination, are bright counts. It was demonstrated that the count rate increases due to the incident light, validating the device’s response through the Néel-Arrhenius law PhysRevB.104.094433. The spintronic Poisson bolometer achieved a noise equivalent temperature difference (NETD) of 103 mK at a 25 Hz frame rate, and high readout speeds of 10 MHz, which is faster than commercial vanadium dioxide detectors. The high sensitivity and the high readout speed make it a promising candidate for state-of-the-art infrared detection.
@@
In other studies in the field, stochastic MTJs have been utilized for random number generation, probabilistic computing, and neuromorphic computing choi2014magnetic; joshi2020mtj; zhang2021recent; maciel2019magnetic; zink2022review. The stochastic magnetization transitions, known as magnetic random telegraph noise (RTN), exhibit relaxation times that follow an exponential distribution, consistent with Poisson event statistics hayakawa2021nanosecond; 2021relaxation. Similarly, detection-based studies have shown exponentially distributed dark-count interarrival times leif_nanolett. Notably, previous studies have examined the statistics of MTJ transitions only in the absence of incident light.
@@
The Poissonian nature of these counts enables time-resolved measurements with high fidelity and superior signal-to-noise ratio (SNR) compared to analog systems under identical conditions. The spintronic Poisson bolometer also exhibits distinct statistics compared to other digital detectors due to the absence of after-pulsing WANG2019202 and its intrinsic Poisson response, irrespective of light statistics. We also investigate the device’s non-probabilistic response. Excitation with a pulsed laser produces current transients, providing insights into heat transfer and transport within the device. This systematic experimental investigation of the response statistics of spintronic Poisson bolometers under illumination advances our understanding of the underlying device physics and guides optimization of sensitivity, response time, dynamic range, and spectral response. Overall, we establish the spintronic Poisson bolometer as a promising platform for probabilistic, high-speed, and high-sensitivity infrared detection at room temperature.
@@ II Results and Discussion
Even without illumination, thermal fluctuations induce spontaneous transitions between the two sensing-layer states, producing dark counts. Illumination generates a hot spot in the transduction layer, which thermalizes through the MTJ stack and increases the probability of transitions in the sensing layer. This results in an increased count rate observed through MTJ resistance readout (Fig. 1(d)). The behavior of these transitions follows the Néel–Arrhenius law for thermally activated switching2021relaxation:
@@
where $\Delta T$ is the temperature change of the detector, $\varepsilon$ is emissivity, $P$ is the incident radiant power, and $A$ is the effective detector area, $G_{\mathrm{th}}$ is the thermal conductance, $C_{\mathrm{th}}$ is the thermal capacitance, and $\omega$ is the angular modulation frequency 8845784.
@@
which reflects the sensitivity of the detector output to temperature change.8845784 For VOx, $k$ is constant over the operating range, whereas for the spintronic Poisson bolometer, the change in count rate follows the Neel-Arrhenius relation leif_nanolett. For small changes in temperature, the count rate increase is
@@
where $dN$ is the increase in counts due to a temperature rise $dT$, $\tau_{0}$ is the relaxation time, $E_{b}$ is the energy barrier, $k_{B}$ is the Boltzmann constant (Supplementary I). Thus, in both conventional analog bolometers and digital spintronic Poisson bolometers, the signal is linearly dependent on temperature.
@@
where $\tau_{\text{diffusion}}$ is the diffusion time, $\tau_{\text{thermal}}$ is the thermal time constant, and $\tau_{\text{intrinsic}}$ represents intrinsic system dynamics as magnetic spin switching. For VOx bolometers, the response time is dominated by thermal time constant (10 ms), whereas for the spintronic Poisson bolometer, it is governed by the diffusion time (2.2-3.5 ns)Abdullah2023; leif_nanolett. Because the spintronic Poisson bolometer is a probabilistic Poisson detector, deterministic readout of a transition is further limited by approximately 4.6 times the relaxation time (0.1-10 $\mu s$) (Supplementary II). Overall, the spintronic Poisson bolometer is faster than $VO_{x}$ by three orders of magnitude, making it a strong candidate for high-speed detector applications.
@@ II.1 Device Response to 808 nm Modulated Laser
Figure  2 (b) shows the device response to the 808 nm modulated laser. Counts are superimposed on the non-probabilistic voltage signal. When the laser is ON, we observe an increase in device resistance due to laser heating of the electrical contacts as well as an increase in count rate. To understand the statistics of ON and OFF regions, we divide each period into 100 bins (100 ns per bin) and the probability of detecting a count in each bin is computed over 6000 periods, as shown in Fig. 2 (c). The dotted line indicates the time light was incident on the device, and the red line is the voltage of the device averaged over 6000 periods. The probability of a count remains constant over both ON/OFF regions, indicating stochastic behavior which is characteristic of Poisson statistics. When the laser is turned ON, we observe that the probability of a count increases 53 ns after the light is incident (Supplementary).
We then separate the data into laser-ON and laser-OFF regions and study the probability distribution of the number of counts $P(N)$ for each region. We divide the data into 100 time bins and calculate the number of counts that occur in each time bin as shown in Fig. 2 (d). We observe that for both laser OFF and laser ON cases, the count distribution follows a Poisson distribution van1992stochastic:

### v3

# Optical Response in Spintronic Poisson Bolometers

@@ Abstract
Analog bolometers based on temperature-dependent phase-transition materials such as vanadium oxide (VOx) and barium titanate (BTO) represent the state of the art in uncooled infrared detectors. Recently, the first room-temperature spintronic Poisson bolometer based on magnetic tunnel junctions (MTJs) was proposed and demonstrated as a promising infrared detector. Unlike conventional bolometers, the spintronic Poisson bolometer operates in a probabilistic regime dominated by Poissonian noise, where the response is governed by resistance fluctuations arising from thermally activated magnetization transitions. Spontaneous transitions between two metastable magnetic states occur even in the absence of incident light, and the transition probability increases under illumination. In this work, we experimentally study the statistical properties of the optical response of the spintronic Poisson bolometer under illumination. We demonstrate that transitions in spintronic Poisson bolometers, both in the absence and presence of light, exhibit Poissonian behavior, with transition rates and interarrival times modulated by incident radiation. Under illumination, we observe a 153% increase in the count rate accompanied by a 70% reduction in interarrival time. These results establish spintronic Poisson bolometers as a promising platform for high-speed, and high-sensitivity infrared detection at room temperature.
@@ I Introduction
Bolometers are thermal detectors in which incident light causes a temperature-induced resistance change that is read out electrically Yadav et al. (2022). Infrared bolometers are essential for emerging applications in thermal imaging, autonomous navigation, astronomy, and time-of-flight sensing Bao et al. (2023); Lahiri et al. (2012); Sivaprakasam et al. (2020); Kaushal and Kaddoum (2017); Anderson et al. (2023); Rodriguez-Gomez et al. (2023); Day et al. (2024). State-of-the-art uncooled bolometers are analog detectors based on temperature-dependent phase transition of materials such as vanadium oxide (VOx) and barium titanate (BTO) Yeh et al. (2020); Scott et al. (2022). Other uncooled bolometers are based on semiconductor and metallic materials with temperature-dependent resistance, including amorphous silicon and platinumWu et al. (2019); Chen et al. (2021); Lien et al. (2023). In VOx bolometers, major noise sources are Johnson noise, thermal noise, generation–recombination and $1/f$ noise Philipose et al. (2023); Ye et al. (2021); Du et al. (2013). Similarly, the dominant noise in barium titanate (BTO) bolometers and graphene bolometers is Johnson noise Hanel (1961); Du et al. (2013); Safaei et al. (2017). All of the uncooled bolometers aforementioned are analog detectors that produce a continuous voltage or current proportional to the incoming signal, and are dominated by Gaussian noise, as shown in Fig. 1(a).
@@
Recently, the first uncooled spintronic Poisson bolometer, called the spintronic ultrafast nanoscale bolometer (SUN bolometer) Bauer et al. (2025) was proposed. Unlike analog bolometers, the spintronic Poisson bolometers convert incoming signals into discrete counting events and are dominated by Poissonian noise, as shown in Fig. 1(b). In spintronic Poisson bolometers, magnetic tunnel junctions (MTJs) are engineered with an energy barrier of the order of thermal infrared energy at room temperature. It exploits switching between two discrete magnetization states by intrinsic thermal fluctuations and nanoscale heating induced by incident radiation. These digital probabilistic transitions in the absence of light are referred to as dark counts, and when under illumination, are bright counts. It was demonstrated that the count rate increases due to the incident light, validating the device’s response through the Néel-Arrhenius law Chalifour et al. (2021). The spintronic Poisson bolometer achieved a noise equivalent temperature difference (NETD) of 103 mK at a 25 Hz frame rate, and high readout speeds of 10 MHz, which is faster than commercial vanadium dioxide detectors. The high sensitivity and the high readout speed make it a promising candidate for state-of-the-art infrared detection.
@@
In other studies in the field, stochastic MTJs have been utilized for random number generation, probabilistic computing, and neuromorphic computing Choi et al. (2014); Joshi et al. (2020); Zhang et al. (2021); Maciel et al. (2019); Zink et al. (2022). The stochastic magnetization transitions, known as magnetic random telegraph noise (RTN), exhibit relaxation times that follow an exponential distribution, consistent with Poisson event statistics Hayakawa et al. (2021); Kanai et al. (2021). Similarly, detection-based studies have shown exponentially distributed dark-count interarrival times Bauer et al. (2025). Notably, previous studies have examined the statistics of MTJ transitions only in the absence of incident light.
@@
The Poissonian nature of these counts enables time-resolved measurements with high fidelity and superior signal-to-noise ratio (SNR) compared to analog systems under identical conditions. The spintronic Poisson bolometer also exhibits distinct statistics compared to other digital detectors due to the absence of after-pulsing Wang et al. (2019) and its intrinsic Poisson response, irrespective of light statistics. We also investigate the device’s non-probabilistic response. Excitation with a pulsed laser produces current transients, providing insights into heat transfer and transport within the device. This systematic experimental investigation of the response statistics of spintronic Poisson bolometers under illumination advances our understanding of the underlying device physics and guides optimization of sensitivity, response time, dynamic range, and spectral response. Overall, we establish the spintronic Poisson bolometer as a promising platform for probabilistic, high-speed, and high-sensitivity infrared detection at room temperature.
@@ II Results and Discussion
Even without illumination, thermal fluctuations induce spontaneous transitions between the two sensing-layer states, producing dark counts. Illumination generates a hot spot in the transduction layer, which thermalizes through the MTJ stack and increases the probability of transitions in the sensing layer. This results in an increased count rate observed through MTJ resistance readout (Fig. 1(d)). The behavior of these transitions follows the Néel–Arrhenius law for thermally activated switchingKanai et al. (2021):
@@
where $\Delta T$ is the temperature change of the detector, $\varepsilon$ is emissivity, $P$ is the incident radiant power, and $A$ is the effective detector area, $G_{\mathrm{th}}$ is the thermal conductance, $C_{\mathrm{th}}$ is the thermal capacitance, and $\omega$ is the angular modulation frequency Kazmierkowski (2019).
@@
which reflects the sensitivity of the detector output to temperature change.Kazmierkowski (2019) For VOx, $k$ is constant over the operating range, whereas for the spintronic Poisson bolometer, the change in count rate follows the Neel-Arrhenius relation Bauer et al. (2025). For small changes in temperature, the count rate increase is
@@
where $dN$ is the increase in counts due to a temperature rise $dT$, $\tau_{0}$ is the relaxation time, $E_{b}$ is the energy barrier, $k_{B}$ is the Boltzmann constant (Supplementary Material I). Thus, in both conventional analog bolometers and digital spintronic Poisson bolometers, the signal is linearly dependent on temperature.
@@
where $\tau_{\text{diffusion}}$ is the diffusion time, $\tau_{\text{thermal}}$ is the thermal time constant, and $\tau_{\text{intrinsic}}$ represents intrinsic system dynamics as magnetic spin switching. For VOx bolometers, the response time is dominated by thermal time constant (10 ms), whereas for the spintronic Poisson bolometer, it is governed by the diffusion time (2.2-3.5 ns)Abdullah et al. (2023); Bauer et al. (2025). Because the spintronic Poisson bolometer is a probabilistic Poisson detector, deterministic readout of a transition is further limited by approximately 4.6 times the relaxation time (0.1-10 $\mu s$) (Supplementary Material II). Overall, the spintronic Poisson bolometer is faster than $VO_{x}$ by three orders of magnitude, making it a strong candidate for high-speed detector applications.
@@ II.1 Device Response to 808 nm Modulated Laser
Figure  2 (b) shows the device response to the 808 nm modulated laser. Counts are superimposed on the non-probabilistic voltage signal. When the laser is ON, we observe an increase in device resistance due to laser heating of the electrical contacts as well as an increase in count rate. To understand the statistics of ON and OFF regions, we divide each period into 100 bins (100 ns per bin) and the probability of detecting a count in each bin is computed over 6000 periods, as shown in Fig. 2 (c). The dotted line indicates the time light was incident on the device, and the red line is the voltage of the device averaged over 6000 periods. The probability of a count remains constant over both ON/OFF regions, indicating stochastic behavior which is characteristic of Poisson statistics. When the laser is turned ON, we observe that the probability of a count increases 53 ns after the light is incident (Supplementary Material).
We then separate the data into laser-ON and laser-OFF regions and study the probability distribution of the number of counts $P(N)$ for each region. We divide the data into 100 time bins and calculate the number of counts that occur in each time bin as shown in Fig. 2 (d). We observe that for both laser OFF and laser ON cases, the count distribution follows a Poisson distribution Van Kampen (1992):
@@ III Conclusion
IV Supplementary Material
The Supplementary Material provides a detailed derivation of equation (5) and the integration time required for a spintronic Poisson bolometer to produce a deterministic response. Additional device characterization results are also included.
@@ Acknowledgements.
Data Availability Statement
Data available on request from the authors.